OpenStudy (anonymous):

Which inequality represents the following sentence? Seven fewer than a number is at least 14 A. x-7<14 B. 7-x≥14 C. x+7≤14 D. x-7≥14 Can you explain this to me rather than just give me the answer? Thank You

OpenStudy (anonymous):

at least means x-7 should be 14 or more. \[x-7\ge 14\]

OpenStudy (anonymous):

I also need help with this one As a salesperson, Hannah earns $1,200 a month plus 7.5% commission on her sales. She needs to earn at least $3,300 a month to pay her bills. This following inequality can be used to find the amount of sales Hanna needs to have each month. 1200 + 0.075x ≥ 3000 What is the solution to the inequality? A. x ≤ 28,000 B. x ≥ 28,000 C. x ≤ 33,750 D. x ≥ 33,750

OpenStudy (anonymous):

Thanks for the help with the first one

OpenStudy (anonymous):

\[1200+0.075 x \ge3300,\frac{ 75 }{1000 }x \ge 3300-1200\] \[x \ge 2100\times \frac{ 1000 }{75 }~or~x \ge 28000 $\]